# Conformal four-point ladder integrals in diverse dimensions and polylogarithms

### v1

@@ Introduction
The case $D=2$ requires special consideration and there are two possible approaches. One is to perform the limit $D\to 2$ ($\lambda\to 0$) in (2.23) using the relation between Gegenbauer and Chebyshev polynomials. We discuss this method in Appendix A. Another, which we adopt here, is to introduce complex coordinates $x={\sf x}_{1}+{\sf i}{\sf x}_{2}\,,\bar{x}={\sf x}_{1}-{\sf i}{\sf x}_{2}$ and $y={\sf y}_{1}+{\sf i}{\sf y}_{2}\,,\bar{y}={\sf y}_{1}-{\sf i}{\sf y}_{2}$, in (2.18). Indeed, for $D=2$ and $n\geq 1$ we have
@@
$\frac{y\bar{y}}{x\bar{x}}=z\bar{z}\ \ ;\ \ \frac{y}{x}+\frac{\bar{y}}{\bar{x}}=z+\bar{z}\qquad\Rightarrow z=\frac{y}{x}\ \ ;\ \ \bar{z}=\frac{\bar{y}}{\bar{x}}$ (2.25)
@@
$\frac{{\sf x}^{\mu_{1}...\mu_{n}}\,{\sf y}^{\mu_{1}...\mu_{n}}}{({\sf x}^{2}\,{\sf y}^{2})^{n/2}}=\frac{1}{2^{n}}\frac{z^{n}+\bar{z}^{n}}{(z\bar{z})^{n/2}}\qquad\text{for }n\geq 1,$ $\frac{{\sf x}^{\mu_{1}...\mu_{n}}\,{\sf y}^{\mu_{1}...\mu_{n}}}{({\sf x}^{2}\,{\sf y}^{2})^{n/2}}\overset{n=0}{=}1.$ (2.26)
In $D=2$ the expression (2.10) for $\mu(n)$ looks simpler
@@
and finally the general formula (2.18) is reduced in $D=2$ to the following form
@@
In Appendix A we show that, by means of the symmetry property $\tau_{n,\nu}(\beta;0)=\tau_{-n,\nu}(\beta;0)$, one can write (2.27) in the concise form (see (A.7))
@@
(these symmetries for $L=1$, $\beta=1$ were exploited in [81]).
Remark 4. It is well known [82] that the $n$-point conformal vertex, for which the sum of indices on the lines satisfies
@@
transforms under conformal inversions $x_{i}\to\frac{1}{x_{i}}$ as follows
@@
It is evident that an integral $I(x_{1},...,x_{n};\vec{\beta})$, which is depicted as a Feynman graph with $n$ external lines having indices $\vec{\beta}=(\beta_{1},...,\beta_{n})$, and with all internal boldface vertices being conformal, transforms under inversions exactly as conformal $n$-point vertex: $I(\frac{1}{x_{1}},...,\frac{1}{x_{n}};\vec{\beta}))=\prod\limits_{i=1}^{n}x_{i}^{2\beta_{i}}I(x_{1},...,x_{n};\vec{\beta}))$. Then, a conformal invariant function can be chosen as
@@
where conformal cross-ratios are $u^{ij}_{kl}=x_{ij}^{2}x_{kl}^{2}/(x_{ik}^{2}x_{jl}^{2})$ and $(i,j,k,l)$ are all possible different 4-element subsets of the $n$ element set (1,2,,n)(1,2,...,n). Formulas (2.4), (2.6) are examples of these general rules.
Remark 5. Note that one can shift the indices on the lines by one unit in the integrals (2.1), (2.2) by making use the generalized integration-by-parts formula for the $n$-point vertex (see [50])
@@
Dimensional shift for generic propagator index $\beta$
The relations of (conformal) Feynman integrals in different dimensions by means of a special operator acting on the variables of the external legs were considered in many papers; see e.g. [44, 83, 10, 84, 76, 73, 74, 75].
Let us redefine (2.16), (2.23) and introduce the function, which depends only on conformal kinematic parameters $z,\bar{z}$:
@@
Define operator [85, 76, 73, 74, 75]
@@
which was indicated in [76, 73, 74, 75] as a shift operator in dimension of the space-time $D\to D+2$ for conformal (ladder) 4-point diagrams (see also [86, Theorem 49]). Indeed, in our special case we have the following statement.
@@
The function $\widetilde{\Phi}^{(\beta)}_{L}(z,\bar{z};\lambda)$, introduced in (2.31), obeys equation
@@
Let us apply raising operator $R_{d}$ (2.35) to the right-hand side of (2.31). Due to $R_{d}\,(z\bar{z})=0$, it acts nontrivially only on the Gegenbauer polynomials and in view of (2.39) the main formula is
@@
First, we note that after differentiating (A.2) with respect to $r$, we deduce (see [87], section 10.11)
@@
By making use this relation and identity $\tau_{n+1,\nu}(\beta;0)=\tau_{n,\nu}(\beta;1)$, we find for the action of $R_{d}$ to (A.5), (2.33):
@@
where we take into account that the first term in the r.h.s. of (A.5), (2.33) is a zero mode of the operator $R_{d}$. ∎
Loop shift for generic propagator index $\beta$
It is known (see [76, 73, 74, 75]) that together with the dimensional shift operator (2.35), there exists an additional operator inducing recursion in number of loops. Remarkably, this operator differs form the Laplace operator, which in the case of diagrams in Fig. 1 also reduces the number of loops (for the reference see [51, 62, 86, 88]). In this subsection we show that combining operator $H_{\beta}$ with the dimensional shift operator we can built an operator which increases number of loops.
In our work [70] we proved the following spectral relations
$H_{\beta}\;|\psi^{\mu_{1}...\mu_{n}}_{\nu,\lambda}\rangle=\tau_{n,\nu}(\beta;\lambda)\;|\psi^{\mu_{1}...\mu_{n}}_{\nu,\lambda}\rangle\;,\qquad\qquad\widehat{H}_{\beta}\;\langle{\sf x}|\psi^{\mu_{1}...\mu_{n}}_{\nu,\lambda}\rangle=\tau_{n,\nu}(\beta;\lambda)\;\langle{\sf x}|\psi^{\mu_{1}...\mu_{n}}_{\nu,\lambda}\rangle,$ (2.44)
@@
$\langle{\sf x}|\psi^{\mu_{1}...\mu_{n}}_{\nu,\lambda}\rangle=\frac{{\sf x}^{\mu_{1}...\mu_{n}}}{(\sf x^{2})^{\frac{\lambda+1+n}{2}+{\sf i}\nu}}\;,\qquad\qquad\tau_{n,\nu}(\beta;\lambda)=4^{\beta}\frac{\Gamma\left(\frac{\lambda+1+n}{2}+\beta-{\sf i}\nu\right)\Gamma\left(\frac{\lambda+1+n}{2}+{\sf i}\nu\right)}{\Gamma\left(\frac{\lambda+1+n}{2}-\beta+{\sf i}\nu\right)\Gamma\left(\frac{\lambda+1+n}{2}-{\sf i}\nu\right)}\;.$
In the two-dimensional case ($\lambda$ = 0), the symmetric traceless tensor ${\sf x}^{\mu_{1}...\mu_{n}}$ has two components, which in terms of complex coordinates (2.24) take the simple form $x^{n}$ and $\bar{x}^{n}$. Then, we have ${\sf x}^{2}=\bar{x}x$ and using complex coordinates write (2.44) as two relations for $n\geq 0$
@@
We substitute $x=z,\bar{x}=\bar{z}$ in the definition of the two-dimensional operator $\widehat{H}_{\beta}$ and introduce operator
@@
Now we formulate the following statement
@@
The function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};\lambda)$, introduced in (2.31), in the case $\lambda\in\mathbb{Z}_{>0}$ obeys the equations
@@
where operator $R_{\ell}^{(\beta)}(\lambda)$ is defined in (2.47), i.e. it translates the expressions for $L$-loop conformal $4$-point ladder integral to the $(L+1)$-loop conformal $4$-point ladder integral.
Proof. First, we represent (2.45) and (2.46) for $\beta\to-\beta$ and $x=z,\bar{x}=\bar{z}$ as follows ($n\geq 0$)
@@
By means of the symmetry (A.6): $\tau_{n\,,\nu}(\beta;0)=\tau_{-n\,,\nu}(\beta;0)$, and after change $\nu\to-\nu$, we write (2.49), (2.50) as one relation for $n\in\mathbb{Z}$
@@
Comparing the left-hand and right-hand sides of this chain of relations for $L\to L+1$ we deduce (2.48). ∎
Let us rewrite the operator $\widehat{H}_{-\beta}=(-i\partial_{\sf x})^{-2\beta}({\sf x})^{-2\beta}$ in two-dimensional case in terms of the complex coordinates (see (2.24))
@@
In order to analyze the operator $\big(R_{\ell}^{(\beta)}(\lambda)\big)^{-1}$, which shifts the loop number in the opposite direction $L\to L-1$ we need the operator
@@
where we used the explicit expression (2.35) for $R_{d}$. Here we also present the simplest example of the inverse relation to the relation (2.48) for $\beta=1$ and $D=4$ ($\lambda=1$):
@@
where we employed the inverse operator (2.53) for $\beta=1$. This relation can be checked explicitly using the representation for the function $\widetilde{\Psi}^{(1)}_{L}(z,\bar{z};1)$ which we derive in subsection 3.1. We provide the corresponding calculations in Appendix B.
Remark 6. In [76] the authors demonstrated that the operator (cf. (2.35))
@@
This relation for $\lambda=1$ and relation (2.54) gives us the linear differential equation on the function $\widetilde{\Phi}^{(1)}_{L}(z,\bar{z};1)$:
@@
Remark 7. Another way of inducing the loop recursions relies on the application of effective two-dimensional Laplace operator. For example, authors of [88] consider $D$-dimensional ladder diagrams, introducing function, which in the case $\lambda=1$ is connected with $\widetilde{\Phi}^{(1)}_{L}(z,\bar{z};1)$ by a rule (see [88, eq. (3.6)])
@@
The step of recursion then consist of using relation [88, eq. (3.2)] together with the following amputation of the line connecting $z$ and 0 (see figure in [88, eq. (3.5)]), which results in
@@
Using rescaling (2.60) in this relation one immediately obtains (2.54).
Ladder integrals for even dimensions and integer $\beta$
It is desirable to evaluate the conformal integrals (2.1) and (2.2), corresponding to the $L$-loop ladder diagrams in Fig. 1, for an arbitrary choice of the parameter $\beta$ (assuming fixing the indices of lines as in (2.7), namely index $\beta$ on the horizontal lines and $D/2-\beta$ on the vertical lines). The one-integral representation (2.23) provides such a possibility, since the integration over $\nu$ can be performed evaluating residues. The analytical result, however, seems to be feasible only for $\beta\in\mathbb{Z}_{>0}$ due to the significant simplifications in the pole structure.
In the following subsection 3.1 we argue that the one-integral representation (2.23) is sufficient to obtain the explicit analytical result in the case of even $D$ and $\beta=1$. We perform such a derivation and show that corresponding representation with Gegenbauer polynomials (2.23) for $\beta=1$ can be systematically rewritten in the form that consists of classical polylogarithms with the coefficients being rational functions in $z$ and $\bar{z}$. The pole structure of (2.23) in the case of $\beta=2,3\dots$ for $D\geq 4$ is quite involved, so instead of direct residue calculations we use the approach based on dimensional shift identities. Thus, in subsection 3.2 we discuss the remarkable factorization occuring in the two-dimensional answer, which with the help of (2.35) can be translated to any even dimension (note that this result holds for arbitrary $\beta$). In subsection 3.3 we then use this result for the case of $\beta=2,3,\ldots$. Of particular interest is the mechanism of regularization of infrared singularities arising for the special combinations of $D$ and $\beta$.
Ladder integrals and polylogarithms for $\beta=1$

### v2

@@ 2.1 Formulation of the problem and introductory remarks
The case $D=2$ requires special consideration and there are two possible approaches. One is to perform the limit $D\to 2$ ($\lambda\to 0$) in (2.23) using the relation between the Gegenbauer and Chebyshev polynomials. We discuss this method in Appendix A. Another thing we adopt here is the introduction of the complex coordinates $x={\sf x}_{1}+{\sf i}{\sf x}_{2}\,,\bar{x}={\sf x}_{1}-{\sf i}{\sf x}_{2}$ and $y={\sf y}_{1}+{\sf i}{\sf y}_{2}\,,\bar{y}={\sf y}_{1}-{\sf i}{\sf y}_{2}$, in (2.18). Indeed, for $D=2$ and $n\geq 1$ we have
@@
$\frac{y\bar{y}}{x\bar{x}}=z\bar{z}\ \ ;\ \ \frac{y}{x}+\frac{\bar{y}}{\bar{x}}=z+\bar{z}\quad\Rightarrow z=\frac{y}{x}\ \ ;\ \ \bar{z}=\frac{\bar{y}}{\bar{x}}$ (2.25)
@@
$\frac{{\sf x}^{\mu_{1}...\mu_{n}}\,{\sf y}^{\mu_{1}...\mu_{n}}}{({\sf x}^{2}\,{\sf y}^{2})^{n/2}}=\frac{1}{2^{n}}\frac{z^{n}+\bar{z}^{n}}{(z\bar{z})^{n/2}}\quad\text{for }n\geq 1,$ $\frac{{\sf x}^{\mu_{1}...\mu_{n}}\,{\sf y}^{\mu_{1}...\mu_{n}}}{({\sf x}^{2}\,{\sf y}^{2})^{n/2}}\overset{n=0}{=}1.$ (2.26)
In $D=2$ expression (2.10) for $\mu(n)$ looks simpler
@@
and finally the general formula (2.18) is reduced in $D=2$ to the following form:
@@
In Appendix A we show that by means of the symmetry property $\tau_{n,\nu}(\beta;0)=\tau_{-n,\nu}(\beta;0)$ one can write (2.27) in the concise form (see (A.7))
@@
(these symmetries for $L=1$, $\beta=1$ were exploited in Loebbert et al. (2020)).
Remark 4. It is well known Symanzik (1972) that the $n$-point conformal vertex, for which the sum of indices on the lines satisfies
@@
transforms under conformal inversions $x_{i}\to\frac{1}{x_{i}}$ as follows:
@@
It is evident that the integral $I(x_{1},...,x_{n};\vec{\beta})$, which is depicted as a Feynman graph with $n$ external lines having indices $\vec{\beta}=(\beta_{1},...,\beta_{n})$ and with all internal boldface vertices being conformal, transforms under inversions exactly as a conformal $n$-point vertex: $I(\frac{1}{x_{1}},...,\frac{1}{x_{n}};\vec{\beta}))=\prod\limits_{i=1}^{n}x_{i}^{2\beta_{i}}I(x_{1},...,x_{n};\vec{\beta}))$. Then, a conformal invariant function can be chosen as
@@
where the conformal cross-ratios are $u^{ij}_{kl}=x_{ij}^{2}x_{kl}^{2}/(x_{ik}^{2}x_{jl}^{2})$ and $(i,j,k,l)$ are all possible different 4-element subsets of the $n$ element set (1,2,,n)(1,2,...,n). Formulas (2.4), (2.6) are examples of these general rules.
Remark 5. Note that one can shift the indices on the lines by one unit in the integrals (2.1), (2.2) by making use of the generalized integration-by-parts formula for the $n$-point vertex (see Gorishnii and Isaev (1985))
@@
2.2 Dimensional shift for generic propagator index $\beta$
The relations of (conformal) Feynman integrals in different dimensions (by means of a special operator acting on the variables of the external legs) were considered in many papers; see e.g. Drummond et al. (2009); Tarasov (1996); Lee (2015); Paulos et al. (2012); Loebbert and Stawinski (2024); Petkou (2021); Karydas et al. (2024, 2025).
Let us redefine (2.16), (2.23) and introduce the function, which depends only on the conformal kinematic parameters $z,\bar{z}$:
@@ 2.2 Dimensional shift for generic propagator index $\beta$
Define the operator Simmons-Duffin (2014); Loebbert and Stawinski (2024); Petkou (2021); Karydas et al. (2024, 2025)
@@
which was indicated in Loebbert and Stawinski (2024); Petkou (2021); Karydas et al. (2024, 2025) as a shift operator in dimension of the space-time $D\to D+2$ for conformal (ladder) 4-point diagrams (see also (Borinsky and Schnetz, 2022a, Theorem 49)). Indeed, in our special case we have the following statement.
@@
The function $\widetilde{\Phi}^{(\beta)}_{L}(z,\bar{z};\lambda)$, introduced in (2.31), obeys the equation
$R_{d}\,\widetilde{\Phi}^{(\beta)}_{L}(z,\bar{z};\lambda)=\widetilde{\Phi}^{(\beta)}_{L}(z,\bar{z};\lambda+1)\;,$ (2.36)
@@
Let us apply the raising operator $R_{d}$ (2.35) to the right-hand side of (2.31). Due to $R_{d}\,(z\bar{z})=0$, it acts nontrivially only on the Gegenbauer polynomials, and in view of (2.39) the main formula is
@@
First, we note that after differentiating (A.2) with respect to $r$, we deduce (see Erdélyi and Bateman (1953b), section 10.11)
@@
By making use of this relation and identity $\tau_{n+1,\nu}(\beta;0)=\tau_{n,\nu}(\beta;1)$, we find for the action of $R_{d}$ on (A.5), (2.33):
@@
where we take into account that the first term in the r.h.s. of (A.5), (2.33) is the zero mode of the operator $R_{d}$. ∎
2.3 Loop shift for generic propagator index $\beta$
It is known (see Loebbert and Stawinski (2024); Petkou (2021); Karydas et al. (2024, 2025)) that together with the dimensional shift operator (2.35) there exists an additional operator inducing recursion in the number of loops. Remarkably, this operator differs form the Laplace operator, which in the case of diagrams in Fig. 1, also reduces the number of loops (for the reference see Drummond et al. (2007); Drummond (2013); Borinsky and Schnetz (2022a, b)). In this subsection we show that combining the operator $H_{\beta}$ with the dimensional shift operator, we can built an operator which increases the number of loops.
In our work Derkachov et al. (2023), we proved the following spectral relations:
$H_{\beta}\;|\psi^{\mu_{1}...\mu_{n}}_{\nu,\lambda}\rangle=\tau_{n,\nu}(\beta;\lambda)\;|\psi^{\mu_{1}...\mu_{n}}_{\nu,\lambda}\rangle\;,\quad\quad\widehat{H}_{\beta}\;\langle{\sf x}|\psi^{\mu_{1}...\mu_{n}}_{\nu,\lambda}\rangle=\tau_{n,\nu}(\beta;\lambda)\;\langle{\sf x}|\psi^{\mu_{1}...\mu_{n}}_{\nu,\lambda}\rangle,$ (2.44)
@@ 2.3 Loop shift for generic propagator index $\beta$
$\langle{\sf x}|\psi^{\mu_{1}...\mu_{n}}_{\nu,\lambda}\rangle=\frac{{\sf x}^{\mu_{1}...\mu_{n}}}{(\sf x^{2})^{\frac{\lambda+1+n}{2}+{\sf i}\nu}}\;,\quad\quad\tau_{n,\nu}(\beta;\lambda)=4^{\beta}\frac{\Gamma\left(\frac{\lambda+1+n}{2}+\beta-{\sf i}\nu\right)\Gamma\left(\frac{\lambda+1+n}{2}+{\sf i}\nu\right)}{\Gamma\left(\frac{\lambda+1+n}{2}-\beta+{\sf i}\nu\right)\Gamma\left(\frac{\lambda+1+n}{2}-{\sf i}\nu\right)}\;.$
In the two-dimensional case ($\lambda$ = 0), the symmetric traceless tensor ${\sf x}^{\mu_{1}...\mu_{n}}$ has two components, which in terms of complex coordinates (2.24) take the simple form $x^{n}$ and $\bar{x}^{n}$. Then, we have ${\sf x}^{2}=\bar{x}x$ and using the complex coordinates write (2.44) as two relations for $n\geq 0$
@@
We substitute $x=z,\bar{x}=\bar{z}$ in the definition of the two-dimensional operator $\widehat{H}_{\beta}$ and introduce the operatorSince $R_{d}$ has zero modes, the operator $(R_{d})^{-\lambda}$ must be understood formally as an operator sending the function $\widetilde{\Phi}^{(\beta)}_{L}(z,\bar{z};\lambda)$ to the function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};0)$, the explicit form of which is known (see Proposition 6 below).
@@
Now we formulate the following statement:
@@
The function $\widetilde{\Phi}_{L}^{(\beta)}(z,\bar{z};\lambda)$ introduced in (2.31) in the case $\lambda\in\mathbb{Z}_{>0}$ obeys the equations
@@
where the operator $R_{\ell}^{(\beta)}(\lambda)$ is defined in (2.47), i.e. it translates the expressions for the $L$-loop conformal $4$-point ladder integral to the $(L+1)$-loop conformal $4$-point ladder integral.
Proof. First, we represent (2.45) and (2.46) for $\beta\to-\beta$ and $x=z,\bar{x}=\bar{z}$ as follows ($n\geq 0$):
@@
By means of the symmetry (A.6): $\tau_{n\,,\nu}(\beta;0)=\tau_{-n\,,\nu}(\beta;0)$, and after the change $\nu\to-\nu$, we write (2.49), (2.50) as one relation for $n\in\mathbb{Z}$
@@
Comparing the left-hand and right-hand sides of this chain of relations for $L\to L+1$, we deduce (2.48). ∎
Let us rewrite the operator $\widehat{H}_{-\beta}=(-i\partial_{\sf x})^{-2\beta}({\sf x})^{-2\beta}$ in the two-dimensional case in terms of the complex coordinates (see (2.24))
@@
In order to analyze the operator $\big(R_{\ell}^{(\beta)}(\lambda)\big)^{-1}$, which shifts the loop number in the opposite direction $L\to L-1$, we need the operator
@@
where we used the explicit expression (2.35) for $R_{d}$. Here we also present the simplest example of the inverse relation to relation (2.48) for $\beta=1$ and $D=4$ ($\lambda=1$):
@@
where we employed the inverse operator (2.53) for $\beta=1$. This relation can be checked explicitly using the representation for the function $\widetilde{\Phi}^{(1)}_{L}(z,\bar{z};1)$ which we derive in subsection 3.1. We provide the corresponding calculations in Appendix B.
Remark 6. In Loebbert and Stawinski (2024), the authors demonstrated that the operator (cf. (2.35))
@@
This relation for $\lambda=1$ and relation (2.54) give us the linear differential equation for the function $\widetilde{\Phi}^{(1)}_{L}(z,\bar{z};1)$:
@@
Remark 7. Another way of inducing the loop recursions relies on the application of the effective two-dimensional Laplace operator. For example, the authors of Borinsky and Schnetz (2022b) considered $D$-dimensional ladder diagrams by introducing a function which, in the case $\lambda=1$, is connected with $\widetilde{\Phi}^{(1)}_{L}(z,\bar{z};1)$ by the rule (see (Borinsky and Schnetz, 2022b, eq. (3.6)))
@@
The step of recursion then consists in using relation (Borinsky and Schnetz, 2022b, eq. (3.2)) together with the following amputation of the line connecting $z$ and 0 (see figure in (Borinsky and Schnetz, 2022b, eq. (3.5))), which results in
@@
Using rescaling (2.60) in this relation, one immediately obtains (2.54).
3 Ladder integrals for even dimensions and integer $\beta$
It is desirable to evaluate the conformal integrals (2.1) and (2.2), corresponding to the $L$-loop ladder diagrams in Fig. 1 for an arbitrary choice of the parameter $\beta$ (assuming fixing the indices of lines as in (2.7), namely, index $\beta$ on the horizontal lines and $D/2-\beta$ on the vertical lines). The one-integral representation (2.23) provides such a possibility, since the integration over $\nu$ can be performed by evaluating residues. The analytical result, however, seems to be feasible only for $\beta\in\mathbb{Z}_{>0}$ due to the significant simplifications in the pole structure.
In the following subsection 3.1, we argue that the one-integral representation (2.23) is sufficient to obtain an explicit analytical result in the case of even $D$ and $\beta=1$. We perform such a derivation and show that the corresponding representation with Gegenbauer polynomials (2.23) for $\beta=1$ can be systematically rewritten in the form that consists of classical polylogarithms with the coefficients being rational functions in $z$ and $\bar{z}$. The pole structure of (2.23) in the case of $\beta=2,3\dots$ for $D\geq 4$ is quite involved, so instead of direct residue calculations we use the approach based on dimensional shift identities. Thus, in subsection 3.2, we discuss the remarkable factorization occurring in the two-dimensional answer which, with the help of (2.35), can be translated to any even dimension (note that this result holds for arbitrary $\beta$). In subsection 3.3, we then use this result for the case of $\beta=2,3,\ldots$. Of particular interest is the mechanism of regularization of infrared singularities arising for the special combinations of $D$ and $\beta$.
3.1 Ladder integrals and polylogarithms for $\beta=1$